Write the equation of the line that satisfies the given conditions. Express the final equation in standard form.
x-intercept is -2 and the y-intercept is 5

Answer by Cromlix(4381)

Hi there,
First gradient = 5/2 = 2.5 (y distance/x distance)
Then you have the y intercept = 5 (c)
y = mx + c
y = 2/5x + 5
Hope this helps :-)